5 years ago, claire was half the age she is in four years. what will his age in six years?
Answer by ankor@dixie-net.com(22740)

5 years ago, claire was half the age she is in four years.
let c = C's age now
c - 5 = .5(c+4)
c - 5 = .5c + 2
c -.5c = 2 + 5
.5c = 7
c = 14 yrs now
:
what will his age in six years?
14 + 6 = 20 yrs